Circular No. 8071

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVA 2003ap Further to IAUC 8067, W. M. Wood-Vasey, G. Aldering, and P. Nugent

report the discovery of an apparent supernova (mag 16.7) on unfiltered NEAT images taken with the Palomar 1.2-m Schmidt telescope on Feb. 9.31, 9.33, and 9.35 UT. SN 2003ap is located at α = 9^{h}50^{m}52^{s}.83, δ = −21^{o}51′10′′.1 (equinox 2000.0), which is 0′′.7 west and 0′′.9 south of the center of the apparent host galaxy. No point source is visible at this location on NEAT images from 2002 Jan. 1, 2, and 2003 Jan. 12 (limiting mag 19.5).

N. Suntzeff, Cerro Tololo Interamerican Observatory (CTIO), reports that a CCD spectrogram (range 570–750 nm) of SN 2003ap, obtained with the CTIO 1.5-m telescope on Feb. 11.20 UT, reveals a supernova spectrum (consistent with a type-Ia supernova roughly near or before maximum) that is superimposed on an active-galaxy spectrum showing a single broad absorption feature that is probably the Si II doublet at 636 nm, blueshifted by 10400 km/s (in the frame of the host galaxy). The galaxy redshift is z = 0.030, based on Hα and [S II] 672-nm nuclear emission.

SUPERNOVAE 2003af AND 2003ag T. Matheson, P. Challis, and R. Kirshner, Harvard-Smithsonian Center

for Astrophysics, report that a spectrum (range 370–750 nm) of SN 2003af (cf. IAUC 8066), obtained by P. Berlind on Feb. 9.37 UT with the F. L. Whipple Observatory 1.5-m telescope (+ FAST spectrograph), shows it to be a type-Ia supernova near maximum. Narrow emission lines superposed on the spectrum of the supernova indicate a recession velocity of 6100 km/s for the host galaxy. Using this value, the supernova expansion velocity is ∼ 12000 km/s for Si II (rest 635.5 nm). The spectral-feature age of the supernova is 1 ± 2 days after maximum light.

Matheson et al. also report that a spectrum of SN 2003ag (cf. IAUC 8067), obtained as above on Feb. 9.39 UT, shows it to be a type-Ia supernova near maximum. Adopting the NASA/IPAC Extragalactic Database recession velocity of 6777 km/s for the host galaxy, the supernova expansion velocity is 11500 km/s for Si II (rest 635.5 nm). The spectral-feature age of the supernova is 3 ± 2 days before maximum light.

Suntzeff adds that a CCD spectrum (range 570–750 nm) of SN 2003ag, obtained at CTIO as above on Feb. 11.25 UT, is consistent with a type-Ia supernova roughly near maximum light, based on the strong absorption feature at 625.5 nm. Associating this with the Si II doublet at 636 nm, the rest-frame expansion velocity of the supernova would be 11400 km/s.

2003 February 12  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2003 CBAT  Daniel W. E. Green